OpenStudy (help!!!!):

Anyone help out with this quick algebra 1 problem. I quite don't understand how to make the equation. I just want the explanation to the answer.

OpenStudy (help!!!!):

A collection of 40 nickels and quarters has a value of $6.40. How many are there of each kind?

OpenStudy (help!!!!):

This is what I got .5x+.25x=40

OpenStudy (help!!!!):

which is wrong...

OpenStudy (anonymous):

x+y=40 0.05x+0.25x=6.40 I'm pretty sure this is the right setup. Then just solve the system of e

OpenStudy (anonymous):

*system of equations.

OpenStudy (anonymous):

oops, for the second equation it should be 0.05x+0.25y. I'm sorry for my typo.

OpenStudy (help!!!!):

opps, could you write in a one step equation

OpenStudy (anonymous):

0.25x+0.1y=6,4

OpenStudy (anonymous):

do u get this

OpenStudy (anonymous):

and then the second one is x+y=40

OpenStudy (anonymous):

Use substitution method

OpenStudy (anonymous):

n = number of nickels 5n = value of the nickels in cents q = number of quarters 25q = value of the quarters in cents . n+q = 40 n = 40-q . 5n + 25q = 640 cents . substitute . 5(40-q) +25q = 640 200 -5q +25q = 640 20q = 440 q = 22 . n = 40-22 n = 18 . check the values to be sure this is the correct answer . 25(22) = 550 cents 5(18) = 90 cents 550 + 90 = 640 cents = $6.40

OpenStudy (anonymous):

x=40-y 0.25(40-y)+y=40

OpenStudy (anonymous):

solve

OpenStudy (anonymous):

Okay, well, you just combine equations real quick. The number of nickels is x, and the number of quarters is (40-x). So write it as 0.05x+0.25(40-x)=6.40. Then solve for x.

OpenStudy (anonymous):

That should help and explain a little bit

OpenStudy (anonymous):

@Brandonsnider188 thats a lot more complicated though

OpenStudy (anonymous):

Not really.

OpenStudy (anonymous):

It's pretty much what you guys are saying but summed up X)

OpenStudy (anonymous):

yes it is mine is 3 steps

OpenStudy (anonymous):

oops didn't notice sry

OpenStudy (anonymous):

Yeah, we both said the same thing in different ways.

OpenStudy (help!!!!):

thanks man :D @Brandonsnider188